if 30000 is 8% of an account what is the total amount of money in the account

Answer by TimothyLamb(4379)

c * 0.08 = 30000
c = 30000 / 0.08 = 375000
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php